# Clothes-Changing Person Re-Identification with Feasibility-Aware Intermediary Matching

### v1

@@ I Introduction
Person re-identification (re-id) [1, 2, 3] aims at identifying the target person across different times and locations captured by surveillance systems. Most existing works [4, 5, 6, 7, 8, 9] assume that a person always wears the same clothes, which may not be applicable in situations where individuals change their clothes. However, clothes changing is commonplace in long-term re-id scenarios. Due to its crucial role in intelligent surveillance systems, clothes-changing re-id has received increasing attention.
@@
In order to address the above challenges, we propose to utilize clothes-relevant features which are often overlooked in clothes-changing scenarios. The clothes-relevant features offer additional advantages over their clothes-irrelevant counterparts. Specifically, clothes-relevant features convey identity information through the clothes appearance. This identity information remains visible and consistent under pose and view changes, ensuring information adequacy and intra-class invariance. Consequently, when a sample lacks clothes-irrelevant clues, by alternatively leveraging clothes-relevant features, we can fetch samples with the same clothes and abundant clothes-irrelevant clues. These samples can then serve as intermediaries for matching clothes-changing targets.
@@
Built upon this inspiration, we employ clothes-relevant features to establish a novel intermediary matching process, which can match samples with insufficient clothes-irrelevant information to more informative intermediate samples. Specifically, as depicted in Fig. 1, for query ($A$) lacking adequate identity information, we can leverage clothes-relevant features to match the query with intermediaries that have richer clothes-irrelevant identity characteristics (e.g., face). For query ($B$) exhibiting large intra-class variation with target ($B$), the clothes-relevant features can be utilized to match the query with intermediaries that have better-aligned clothes-irrelevant clues (e.g., body shape) to the target. This intermediary matching approach streamlines the matching between the query and target individuals, utilizing intermediaries with highly informative clothes-irrelevant characteristics. Notably, given that images of pedestrians wearing the same clothes are easily accessible from nearby frames or multiple camera views in short-term video surveillance, the fundamental feasibility of the intermediary matching process is guaranteed.
Nevertheless, the re-id performance may still be impaired if the desired intermediaries (e.g., same-clothes samples) are unavailable, or if the retrieved intermediaries contain unreliable clothes-irrelevant identity information. Therefore, it becomes imperative to assess the feasibility of intermediary matching by evaluating the quality of intermediaries, considering both their availability and the reliability of their clothes-irrelevant identity information: (1) Availability refers to the accessibility of the desired intermediary samples. Specifically, when matching with clothes-relevant features, availability refers to the accessibility of same-clothes intermediaries. For example, as shown on the left part of Fig. 2(a), if same-clothes samples do not exist in the data source of intermediaries, the availability of intermediaries is low. On the other hand, when matching with clothes-irrelevant features, availability refers to the accessbility of same-identity samples. As shown on the right part of Fig. 2(a), if same-identity samples do not exist in the data source, the availability of intermediaries is low as well. (2) Reliability refers to the integrity of clothes-irrelevant identity cues in the intermediary samples. For example, as shown in Fig. 2(b), intermediaries with clear facial views and full body views possess high reliability, whereas intermediaries with obscured face or occluded body parts have low reliability.
@@ II-A Clothes-changing person re-id
Current clothes-changing person re-id methods [10, 11, 22, 15, 23, 13] focus on extracting clothes-irrelevant identity features. A line of works resorts to multi-modality data, such as contour sketch [20, 24], human silhouette [11, 16], radio signal [25], body keypoints [14] and 3D shape [17, 18]. For instance, the work [14] develops a clothes-elimination shape-distillation framework to extract clothes-irrelevant representations under the guidance of body keypoint embeddings, and the work [16] disentangles clothes-irrelevant cues from clothes-relevant cues by letting the model to reconstruct the silhouettes of clothes-irrelevant and clothes-relevant parts separately. Another type of method extract clothes-irrelevant features using single-modality image data. For example, the work [10] relies on adversarial learning to mine clothes-irrelevant information, the work [13] conducts a clothes-change ID-unchange feature augmentation to boost the decomposition of clothes-independent identity information, and the work [19] proposes an association-forgetting method to facilitate the learning of identity-relevant features through association learning, while precluding the impact of identity-irrelevant features through forgetting learning. All the above works assume that the clothes-irrelevant feature alone is sufficient to represent pedestrians. However, the large variations in human pose and camera view can cause clothes-irrelevant features to lack identity clues and exhibit large intra-class variations. In this work, we propose an intermediary-assisted matching strategy, which additionally utilizes clothes-relevant characteristic to reduce identification difficulty.
@@ II-C Reliability modeling in re-id
In re-id task, a group of works [34, 35, 36, 37, 38] model the reliability of data samples to mitigate the impact of outliers or noisy labels. Recent works [36, 37] predict the sample reliability by mapping each sample to a Gaussian distribution in latent feature space, and learning the variance which represents sample reliability. These works only model the reliability for same-clothes scenarios, and the predicted reliability is only used for training. Differently, our method builds the reliability for clothes-changing scenarios, while managing to apply the reliability score for the testing procedure.
@@ III-A Feature Decoupling Module
where $C_{\varphi}^{C}\left(\cdot\right)$ denotes the clothes classifier with parameter $\varphi$. Notably, for triplet loss $\ell_{tri}\left(\boldsymbol{f}^{re}\right|y^{C})$, we only select negative samples within the same identity category to better distinguish different clothes of the same identity, thereby more effectively mining fine-grained clothes information.
@@ III-B Identity Information Reliability Module
where $\lambda_{fv}$ is the margin to bound the average variance level. By employing $\mathcal{L}_{fv}$, to achieve a more significant reduction in classification loss, the model will prefer to reduce the variance of samples of high reliability over those low-reliability samples. Thus we can ensure that samples with higher reliability exhibit low feature variance. In this way, we can predict the reliability score of clothes-irrelevant identity information as $r^{id}=1.0-\rho$.
@@ III-D Intermediary-Based Feasibility Weighting Module
where $s^{re}$ and $s^{ir}$ denotes the cosine similarity (scaled to $\left[0,1\right]$) of $\boldsymbol{f}^{re}$ and $\boldsymbol{f}^{ir}$, respectively. Similarly, we can compute $s_{\mathcal{B}}$ and $s_{\mathcal{C}}$:
@@ IV-C Comparison with State-of-the-art Methods
TABLE III: Ablation on the effectiveness of IBFW. ‘w/o SC’ denotes same-clothes samples for intermediary matching are excluded. ‘-50% high reliability’ denotes that the samples with top 50% highest reliability are excluded for intermediary matching. ‘all’ denotes no samples are excluded. ‘w/o IM’ denotes conventional re-ranking [32] without our proposed IM. Top-1 and mAP of the clothes-changing (CC) setting is reported.
@@ IV-D Ablation Study
Figure 4: The top-1 accuracy and mAP with different $\lambda_{A}$, $\lambda_{B}$ and $\lambda_{C}$ on the clothes-changing setting of LTCC. In (a) , we fix $\lambda_{B}$ to $0.6$ and $\lambda_{C}$ to $0.1$. In (b) , we fix $\lambda_{A}$ to $0.3$ and $\lambda_{C}$ to $0.1$. In (c) , we fix $\lambda_{A}$ to $0.3$ and $\lambda_{B}$ to $0.6$.
The effectiveness of Intermediary-Based Feasibility Re-weighting (IBFW) module.  In Tab. III, we conduct experiments to validate the effects of our IBFW module. Since IBFW is primarily designed to deal with situations where high-quality intermediaries are inaccessible, we manually simulate these situations and test the performance of IBFW: (1) To simulate situations of low intermediary availability, we create ‘w/o SC’ setting by excluding same-clothes samples when fetching intermediaries with clothes-relevant features. (2) To simulate situations of low clothes-irrelevant identity information reliability, we create ‘-50% high reliability’ setting by excluding samples of the top 50% highest reliability score $r^{id}$ from the gallery. As shown in Tab. III, under both situations, FAIM (‘w/ IBFW’) performs better than not using IBFW (‘w/o IBFW’). By employing IBFW, the performance is less impaired in low availability or low reliability situations, and the advantage against conventional re-ranking methods (‘w/o IM’) can be maintained. Conclusively, IBFW is useful in addressing low-quality intermediaries and guaranteeing the robustness of IM against adverse data conditions.
TABLE IV: Ablation study on the effectiveness of IIR module. ‘w/o IIR denotes training the model without adopting the IIR module, and testing without incorporating the reliability score. Top-1 and mAP of the clothes-changing (CC) setting is reported.
@@
TABLE V: Ablation study on the construction method of $\boldsymbol{\Sigma}^{ir}$ in the Identity Information Reliability block. top-1 and mAP of the clothes-changing (CC) setting on LTCC is reported.
@@
The effectiveness of Identity Information Reliability (IIR) module.  In Tab. IV, we ablate on the effectiveness of our proposed IIR module. As the results in Tab. IV shows, the adoption of IIR brings a top-1 performance gain of $2.5\%$ and $3.3\%$ on LTCC and PRCC datasets, respectively. This verifies the importance of modeling the reliability of clothes-irrelevant information in our FAIM framework.
@@
Furthermore, results in Tab. V support the aforementioned claim. As seen, the final performance when using clothes-changing variance $\boldsymbol{\Sigma}^{c}$ to construct $\boldsymbol{\Sigma}^{ir}$ is better than directly predicting the variance following [36] which is unaware of clothes changes. Moreover, as shown in Tab. V, our approach of constructing $\boldsymbol{\Sigma}^{c}$ in instance-wise manner is advantageous compared to using a global $\boldsymbol{\Sigma}^{c}$ [13] for every instance. The main reason is that different samples exhibit variations in their clothes-changing semantic directions, so that instance-specific clothes-changing variance could model the clothes-changing directions more precisely.
Figure 6: T-SNE [49] visualization results clothes-relevant feature and clothes-irrelevant feature derived from the FD module. Each color represents one identity, and each marker style stands for each type of clothing within each identity.
TABLE VI: Ablation on the incorporation design of clothes-relevant and clothes-irrelevant features. We try to incorporate $f^{ir}$ and $f^{re}$ by concatenation along the channel dimension. ‘Baseline + Concat.’ denotes using concatenated features to perform regular re-ranking [32]. ‘FAIM + Concat.’ denotes using concatenated features to substitute $f^{ir}$ and $f^{re}$, and perform the intermediary matching process. For ‘general’, ‘SC’ and ‘CC’ settings, we report the Top-1 accuracy.
@@
Incorporation design of clothes-relevant and clothes-irrelevant features.  In FAIM, we design an intermediary matching pipeline which leverages clothes-relevant features $f^{re}$ and clothes-irrelevant features $f^{ir}$ to perform indirect matching routes. As defined in Eq. 7, in every matching step, we selectively use $f^{ir}$ or $f^{re}$, but never jointly use both features in one step. To validate the advantage of our selective feature incorporation approach, we compare FAIM against jointly use both features in every matching step. When applying the joint usage approach, we substitute every feature in Eq. 7 with the concatenation of $f^{ir}$ and $f^{re}$. Results in Tab. VI shows that using the concatenation $f^{ir}$ and $f^{re}$ as the feature representation in FAIM (FAIM + Concat.) degrades the performance of FAIM on both LTCC and PRCC datasets, and the superiority of our selective feature incorporation design in FAIM is verified. Moreover, FAIM largely outcompetes the approach of concatenating $f^{ir}$ and $f^{re}$ and use the incorporated feature to perform direct matching (Baseline + Concat.), which proves that our carefully designed intermediary matching routes provide a more proper way to jointly utilize both clothes-relevant and clothes-irrelevant features.
Hyper-parameter analysis.  In our training loss function $\mathcal{L}$ (see Eq. 6), there are two hyper-parameters, $\alpha^{ir}$ and $\alpha^{re}$. We set both of them as $0.5$. Here we conduct an hyper-parameter analysis on these two coefficients. The results are shown in Tab. VII. Under our selection of $\alpha^{ir}$ and $\alpha^{re}$, the model gains the overall best performance on both LTCC and PRCC benchmarks.
@@
TABLE VII: Hyper-parameter analysis. To verify the optimality of our hyper-parameter selection (setting both $\alpha^{ir}$ and $\alpha^{re}$ at $0.5$), we fix $\alpha^{re}$ at $0.5$ and change $\alpha^{ir}$, as well as fixing $\alpha^{re}$ at $0.5$ and change $\alpha^{re}$. We report the Top-1 and mAP results of clothes-changing (CC) setting on all datasets.

### v2

@@ I Introduction
Person re-identification (re-id) [2, 1, 3] aims at identifying the target person across different times and locations captured by surveillance systems. Most existing works [6, 4, 5, 7, 46, 45, 66] assume that a person always wears the same clothes, which may not be applicable in situations where individuals change their clothes. However, clothes changing is commonplace in long-term re-id scenarios. Due to its crucial role in intelligent surveillance systems, clothes-changing re-id has received increasing attention.
@@
In order to address the above challenges, we propose to utilize clothes-relevant features which are often overlooked in clothes-changing scenarios. Though clothes traits might fail to discriminate between different identities that wear similar clothes, we can circumvent this problem by learning clothes-relevant features that contain identity-discriminative information. To supervise the clothes-relevant features, we use specially designed clothes label, where each identity category is further divided into multiple fine-grained categories of clothes. This design ensures that samples with different identities never share the same label, even if they wear same or similar clothes. Supervised by this specially designed clothes label, our clothes-relevant features are capable of discriminating between identities, which suits well to the clothes-changing setting. Moreover, the clothes-relevant features offer additional advantages over their clothes-irrelevant counterparts. Specifically, clothes-relevant features convey identity information through the clothes appearance. This identity information remains visible and consistent under pose and view changes, ensuring information adequacy and intra-class invariance. Consequently, when a sample lacks clothes-irrelevant clues, by alternatively leveraging clothes-relevant features, we can fetch samples with the same clothes and abundant clothes-irrelevant clues. These samples can then serve as intermediaries for matching clothes-changing targets.
@@
Built upon this inspiration, we employ clothes-relevant features to establish a novel intermediary matching process. We introduce the intermediary sample, which refers to an extra data sample that serves as a median between the matching of query and target. In this way, instead of directly matching samples with insufficient clothes-irrelevant information to the target, we can first match them to intermediary samples that are more informative, and then match these intermediaries to target. Specifically, as depicted in Fig. 1, for query ($A$) lacking adequate identity information, we can leverage clothes-relevant features to match the query with intermediaries that have richer clothes-irrelevant identity characteristics (e.g., face). For query ($B$) exhibiting large intra-class variation with target ($B$), the clothes-relevant features can be utilized to match the query with intermediaries that have better-aligned clothes-irrelevant clues (e.g., body shape) to the target. This intermediary matching approach streamlines the matching between the query and target individuals, utilizing intermediaries with highly informative clothes-irrelevant characteristics. Notably, given that images of pedestrians wearing the same clothes are easily accessible from nearby frames or multiple camera views in short-term video surveillance, the fundamental feasibility of the intermediary matching process is guaranteed.
Nevertheless, the intermediary matching process would be less effective if the desired intermediaries (e.g., same-clothes samples) are unavailable, or if the retrieved intermediaries contain unreliable clothes-irrelevant identity information. Under these conditions, the re-id performance may still be impaired. Therefore, it becomes imperative to assess the feasibility of intermediary matching, which refers to how well the matching process could be effectively performed with the fetched intermediaries. To judge the feasibility of intermediary matching process, we propose to evaluate the quality of intermediaries, by considering both their availability and the reliability of their clothes-irrelevant identity information: (1) Availability refers to the accessibility of the desired intermediary samples. Specifically, when matching with clothes-relevant features, availability refers to the accessibility of same-clothes intermediaries. For example, as shown on the left part of Fig. 2(a), if same-clothes samples do not exist in the data source of intermediaries, the availability of intermediaries is low. On the other hand, when matching with clothes-irrelevant features, availability refers to the accessbility of same-identity samples. As shown on the right part of Fig. 2(a), if same-identity samples do not exist in the data source, the availability of intermediaries is low as well. (2) Reliability refers to the integrity of clothes-irrelevant identity cues in the intermediary samples. For example, as shown in Fig. 2(b), intermediaries with clear facial views and full body views possess high reliability, whereas intermediaries with obscured face or occluded body parts have low reliability.
@@ II-A Clothes-changing person re-id
In the original person re-identification task, it is assumed that people do not change their clothes across all time periods. Based on this traditional setting, a series of more challenging tasks are derived. Visible-infrared re-id aims at retrieving the same identity across RGB and infrared modalities.  [53] proposed a modality-specific memory network to learn a more unified feature for cross-modality retrieval. Occluded re-id attempts to tackle situations where people are only partly visible.  [20] treated this problem by designing a part-aware transformer to learn representative part features for identities. Moreover, [54] focuses on cross-resolution re-id, which aims at matching person images with varying resolutions. This work addresses this task by learning a resolution-adaptive representation.  [55] studies the problem of unsupervised re-id by raising a self-similarity learning approach that learns discriminative features by pseudo pairs.
Among these challenging settings, clothes-changing person re-id aims at retrieving same-identity person when they expose to clothes changes. Current clothes-changing person re-id methods [13, 9, 14, 12, 16, 36] focus on extracting clothes-irrelevant identity features. A line of works resorts to multi-modality data, such as contour sketch [8, 49], human silhouette [9, 50, 57], radio signal [18], body keypoints [10] and 3D shape [19, 47]. For instance, the work [10] develops a clothes-elimination shape-distillation framework to extract clothes-irrelevant representations under the guidance of body keypoint embeddings, and the work [50] disentangles clothes-irrelevant cues from clothes-relevant cues by letting the model to reconstruct the silhouettes of clothes-irrelevant and clothes-relevant parts separately. Another line of work leverages facial information [58, 59] to aid re-id in clothes-changing scenarios. Thanks to the rapid development in face detection and recognition techniques [60, 61, 62, 63], the works [64, 65] detect and extract facial representations to serve as a strongly reliable identity information for retrieval. Another type of method extract clothes-irrelevant features using single-modality image data. For example, the work [13] relies on adversarial learning to mine clothes-irrelevant information, the work [36] conducts a clothes-change ID-unchange feature augmentation to boost the decomposition of clothes-independent identity information, and the work [48] proposes an association-forgetting method to facilitate the learning of identity-relevant features through association learning, while precluding the impact of identity-irrelevant features through forgetting learning. All the above works assume that the clothes-irrelevant feature alone is sufficient to represent pedestrians. However, the large variations in human pose and camera view can cause clothes-irrelevant features to lack identity clues and exhibit large intra-class variations. In this work, we propose an intermediary-assisted matching strategy, which additionally utilizes clothes-relevant characteristic to reduce identification difficulty.
@@ II-C Reliability modeling in re-id
In re-id task, a group of works [42, 43, 39, 38, 44] model the reliability of data samples to mitigate the impact of outliers or noisy labels. In weakly-supervised setting, [56] proposes to utilize reliability-aware methodology to estimate the uncertainty of self-generated pseudo labels. Recent works [39, 38] predict the sample reliability by mapping each sample to a Gaussian distribution in latent feature space, and learning the variance which represents sample reliability. These works only model the reliability for same-clothes scenarios, and the predicted reliability is only used for training. Differently, our method builds the reliability for clothes-changing scenarios, while managing to apply the reliability score for the testing procedure.
@@ III-A Feature Decoupling Module
where $C_{\varphi}^{C}\left(\cdot\right)$ denotes the clothes classifier with parameter $\varphi$. Notably, in existing clothes-changing re-ID datasets, since clothes labels are fine-grained identity labels, samples from different identities have different clothes labels even if they actually wear same clothes. Therefore, to more effectively mine fine-grained clothes information in $\ell_{tri}\left(\boldsymbol{f}^{re}\right|y^{C})$, we only select negative samples within the same identity category as the anchor sample, to guarantee that their clothes are not same. Specifically, the formulation of $\ell_{tri}\left(\boldsymbol{f}^{re}\right|y^{C})$ is as follows:
$\ell_{tri}\left(\boldsymbol{f}^{re}|y^{C}\right)=\frac{1}{B}$ $\sum_{a}\left[d(\boldsymbol{f}^{re}_{a},\boldsymbol{f}^{re}_{p})-d(\boldsymbol{f}^{re}_{a},\boldsymbol{f}^{re}_{n})+m\right]_{+}$
$p=\arg\max_{i}$ $\,d(\boldsymbol{f}^{re}_{a},\boldsymbol{f}^{re}_{i}),s.t.\,y^{C}_{i}=y^{C}_{a}.$
$n=\arg\min_{i}$ $\,d(\boldsymbol{f}^{re}_{a},\boldsymbol{f}^{re}_{i}),s.t.\,y^{C}_{i}\neq y^{C}_{a},\quad y^{ID}_{i}=y^{ID}_{a}.$ (3)
where $\{a,p,n\}$ denotes a triplet from a mini-batch with batch size $B$: $a$ is anchor sample, $p$ is the positive sample with the largest distance from $a$, and $n$ is the negative sample with smallest distance from $a$. $d$ is cosine distance metric. $\left[\cdot\right]_{+}$ denotes $\max(\cdot,0)$. $y^{C}$ denotes the clothes label and $y^{ID}$ denotes the identity label.
@@ III-B Identity Information Reliability Module
where $\lambda_{fv}$ is the margin to bound the average variance level. By employing $L_{fv}$, the model will have to find a balance between reducing the classification loss and maintaining a total variance level. This will lead the model to reduce the variance of high-reliability samples and maintain the variance of low-reliability samples. Thus we can ensure that samples with higher reliability exhibit low feature variance. In this way, we can predict the reliability score of clothes-irrelevant identity information as $r^{id}=1.0-\rho$.
@@ III-D Intermediary-Based Feasibility Weighting Module
where $s^{re}$ and $s^{ir}$ denotes the cosine similarity (scaled to $\left[0,1\right]$) of $\boldsymbol{f}^{re}$ and $\boldsymbol{f}^{ir}$, respectively. $r^{id}$ denotes the reliability score computed by the IIR module. Similarly, we can compute $s_{\mathcal{B}}$ and $s_{\mathcal{C}}$:
@@ IV-C Comparison with State-of-the-art Methods
TABLE III: Ablation study on the effectiveness of each matching route ($A,B,C$) in IM module. Top-1 and mAP of the clothes-changing (CC) setting is reported.
A B C LTCC PRCC
top-1 mAP top-1 mAP
43.6 23.1 55.3 57.5
45.7 26.1 57.6 61.1
47.2 26.0 58.8 60.8
44.6 23.2 59.4 60.1
47.0 27.2 58.3 61.7
46.2 26.1 59.6 61.7
48.2 25.6 58.8 61.1
48.2 27.5 60.9 62.0
TABLE IV: Ablation on the effectiveness of IBFW. ‘w/o SC’ denotes same-clothes samples for intermediary matching are excluded. ‘-50% high reliability’ denotes that the samples with top 50% highest reliability are excluded for intermediary matching. ‘all’ denotes no samples are excluded. ‘w/o IM’ denotes conventional re-ranking [27] without our proposed IM. Top-1 and mAP of the clothes-changing (CC) setting is reported.
@@ IV-D Ablation Study
It is also worth noting that in Fig. 4(c), when not applying the IBFW, the performance of route $\mathcal{C}$ declines as the weight increases. This phenomenon can be explained as follows: Compared to route $\mathcal{A}$ and $\mathcal{B}$ which only involve one intermediary, route $\mathcal{C}$ involves two intermediaries. This increases the chance of introducing intermediaries with low availability or reliability, which downgrade the re-id accuracy. When not applying IBFW, the feasibility score $s_{\mathcal{C}}$ of is fixed. Therefore, when encountered with intermediaries with low availability or low reliability, route $s_{\mathcal{C}}$ cannot be down-weighted accordingly to avoid the negative influence of low-quality intermediaries. As a result, route $s_{\mathcal{C}}$ exhibits performance decrease in Fig. 4(c).
The effectiveness of matching routes in Intermediary Matching (IM) module.  In IM module, we designed three matching routes $s_{\mathcal{A}}\sim s_{\mathcal{C}}$ to address different circumstances that need intermediaries to facilitate matching. Tab. III conducts ablations to validate the effectiveness of each route. As shown in Tab. III, (1) Compared to the method not leveraging intermediary matching (the first row), utilizing any of the three routes individually could lead to performance gain. (2) Compared to using all three routes, only deploying one or two routes will cause performance drop. These results show the effectiveness of each matching route, and consolidate the importance of incorporating three matching routes altogether for the IM process.
Figure 4: The top-1 accuracy and mAP with different $s_{A}$, $s_{B}$ and $s_{C}$ on the clothes-changing setting of LTCC. In (a) , we fix $s_{B}$ to $0.6$ and $s_{C}$ to $0.1$. In (b) , we fix $s_{A}$ to $0.3$ and $sa_{C}$ to $0.1$. In (c) , we fix $s_{A}$ to $0.3$ and $s_{B}$ to $0.6$.
The effectiveness of Intermediary-Based Feasibility Re-weighting (IBFW) module.  In Tab. IV, we conduct experiments to validate the effects of our IBFW module. Since IBFW is primarily designed to deal with situations where high-quality intermediaries are inaccessible, we manually simulate these situations and test the performance of IBFW: (1) To simulate situations of low intermediary availability, we create ‘w/o SC’ setting by excluding same-clothes samples when fetching intermediaries with clothes-relevant features. (2) To simulate situations of low clothes-irrelevant identity information reliability, we create ‘-50% high reliability’ setting by excluding samples of the top 50% highest reliability score $r^{id}$ from the gallery. As shown in Tab. IV, under both situations, FAIM (‘w/ IBFW’) performs better than not using IBFW (‘w/o IBFW’). By employing IBFW, the performance is less impaired in low availability or low reliability situations, and the advantage against conventional re-ranking methods (‘w/o IM’) can be maintained. Conclusively, IBFW is useful in addressing low-quality intermediaries and guaranteeing the robustness of IM against adverse data conditions.
TABLE V: Ablation study on the effectiveness of IIR module. ‘w/o IIR denotes training the model without adopting the IIR module, and testing without incorporating the reliability score. Top-1 and mAP of the clothes-changing (CC) setting is reported.
@@
TABLE VI: Ablation study on the construction method of $\boldsymbol{\Sigma}^{ir}$ in the Identity Information Reliability block. top-1 and mAP of the clothes-changing (CC) setting on LTCC is reported.
@@
The effectiveness of Identity Information Reliability (IIR) module.  In Tab. V, we ablate on the effectiveness of our proposed IIR module. As the results in Tab. V shows, the adoption of IIR brings a top-1 performance gain of $2.5\%$ and $3.3\%$ on LTCC and PRCC datasets, respectively. This verifies the importance of modeling the reliability of clothes-irrelevant information in our FAIM framework.
@@
Furthermore, results in Tab. VI support the aforementioned claim. As seen, the final performance when using clothes-changing variance $\boldsymbol{\Sigma}^{c}$ to construct $\boldsymbol{\Sigma}^{ir}$ is better than directly predicting the variance following [39] which is unaware of clothes changes. Moreover, as shown in Tab. VI, our approach of constructing $\boldsymbol{\Sigma}^{c}$ in instance-wise manner is advantageous compared to using a global $\boldsymbol{\Sigma}^{c}$ [36] for every instance. The main reason is that different samples exhibit variations in their clothes-changing semantic directions, so that instance-specific clothes-changing variance could model the clothes-changing directions more precisely.
Figure 6: T-SNE [41] visualization results clothes-relevant feature and clothes-irrelevant feature derived from the FD module. Each color represents one identity, and each marker style stands for each type of clothing within each identity. The identity index and clothes index within each identity are labeled in the graph.
TABLE VII: Ablation on the incorporation design of clothes-relevant and clothes-irrelevant features. We try to incorporate $f^{ir}$ and $f^{re}$ by concatenation along the channel dimension. ‘Baseline + Concat.’ denotes using concatenated features to perform regular re-ranking [27]. ‘FAIM + Concat.’ denotes using concatenated features to substitute $f^{ir}$ and $f^{re}$, and perform the intermediary matching process. For ‘general’, ‘SC’ and ‘CC’ settings, we report the Top-1 accuracy.
@@
Incorporation design of clothes-relevant and clothes-irrelevant features.  In FAIM, we design an intermediary matching pipeline which leverages clothes-relevant features $f^{re}$ and clothes-irrelevant features $f^{ir}$ to perform indirect matching routes. As defined in Eq. 8, in every matching step, we selectively use $f^{ir}$ or $f^{re}$, but never jointly use both features in one step. To validate the advantage of our selective feature incorporation approach, we compare FAIM against jointly use both features in every matching step. When applying the joint usage approach, we substitute every feature in Eq. 8 with the concatenation of $f^{ir}$ and $f^{re}$. Results in Tab. VII shows that using the concatenation $f^{ir}$ and $f^{re}$ as the feature representation in FAIM (FAIM + Concat.) degrades the performance of FAIM on both LTCC and PRCC datasets, and the superiority of our selective feature incorporation design in FAIM is verified. Moreover, FAIM largely outcompetes the approach of concatenating $f^{ir}$ and $f^{re}$ and use the incorporated feature to perform direct matching (Baseline + Concat.), which proves that our carefully designed intermediary matching routes provide a more proper way to jointly utilize both clothes-relevant and clothes-irrelevant features.
Effects of Introducing Clothes-relevant Features.  In the Intermediary Matching (IM) process, we introduce clothes-relevant features to help match more informative intermediaries. To ensure that clothes-relevant features have enough discrimination between different identities in the IM process, in Tab. VIII, we calculate the portion of same-identity samples in the intermediaries when matching with clothes-relevant features, and compared with the method of using original feature $\boldsymbol{f}^{o}$ in the entire IM process (denoted as w/o $\boldsymbol{f}^{re}$). As shown in Tab. VIII, in FAIM method (and each of the matching routes $\mathcal{A}\sim\mathcal{C}$), the portion of same-identity intermediaries when matching with clothes-relevant features are comparable with using the original re-id feature. Moreover, we calculated the average feasibility score of same-identity intermediaries ($s_{Pos}$) and different-identity $s_{Neg}$ intermediaries, respectively. From Tab. VIII, we can observe that the feasibility scores of same-identity samples are higher than those of different-identity samples. Therefore, we can conclude that our clothes-relevant feature representation itself could discriminate different identities, thanks to the identity-discriminative supervision signals. Meanwhile, our IBFW module can further eliminate the effects of different-identity samples, by assigning low feasibility scores to them, thereby decreasing the weights of these samples in the whole matching process.
TABLE VIII: Ablation study on the effects of clothes-relevant features in IM process. ‘Pos.ID(%)’ stands for the portion of intermediaries with the same ID as query, when matching with clothes-relevant feature. ‘$s_{Pos}$’ and ‘$s_{Neg}$’ indicates the average feasibility score of same-identity and different-identity intermediaries, respectively. ‘w/o $\boldsymbol{f}^{re}$’ stands for using original feature $\boldsymbol{f}^{o}$ instead of clothes-relevant feature $\boldsymbol{f}^{re}$ for the entire IM process.
method LTCC PRCC
Pos.ID(%) $s_{Pos}$ $s_{Neg}$ Pos.ID(%) $s_{Pos}$ $s_{Neg}$
w/o $\boldsymbol{f}^{re}$ 79.2 - - 62.1 - -
Route $\mathcal{A}$ 82.6 0.50 0.09 61.6 0.48 0.26
Route $\mathcal{B}$ 79.2 0.47 0.12 61.8 0.48 0.28
Route $\mathcal{C}$ 78.2 0.24 0.07 58.4 0.22 0.15
FAIM 80.0 0.40 0.09 60.6 0.39 0.23
Hyper-parameter analysis.  In our training loss function $\mathcal{L}$ (see Eq. 7), there are two hyper-parameters, $\alpha^{ir}$ and $\alpha^{re}$. We set both of them as $0.5$. Here we conduct an hyper-parameter analysis on these two coefficients. The results are shown in Tab. IX. Under our selection of $\alpha^{ir}$ and $\alpha^{re}$, the model gains the overall best performance on both LTCC and PRCC benchmarks.
@@
TABLE IX: Hyper-parameter analysis. To verify the optimality of our hyper-parameter selection (setting both $\alpha^{ir}$ and $\alpha^{re}$ at $0.5$), we fix $\alpha^{re}$ at $0.5$ and change $\alpha^{ir}$, as well as fixing $\alpha^{re}$ at $0.5$ and change $\alpha^{re}$. We report the Top-1 and mAP results of clothes-changing (CC) setting on all datasets.